OpenStudy (goformit100):

A clock is started at noon. By 10 minutes past 5, the hour hand has turned through: A. 145º B. 150º C. 155º D. 160º

OpenStudy (goformit100):

@Noemi95

OpenStudy (anonymous):

draw it out my friend for additional aid

OpenStudy (anonymous):

Angle traced by hour hand in 12 hrs = 360º Angle traced by hour hand in 5 hrs 10 min: 31/6 hrs= (360/12) * (31/6)= ?

OpenStudy (promisespromises):

You've got it right with what you have put down

OpenStudy (promisespromises):

360/12 = 30 degrees per hour. So 5 *30 = 150 degrees. Then 10 mins is 1/6 of an hour and 1/6 x 30 = 5 degrees. And 150+5=155 degrees.